how many liters of chlorine will be needed to make 7.50 moles of C2H2Cl4 at STP
1
Expert's answer
2020-12-08T05:32:30-0500
"C_2H_2 + 2Cl_2 = C_2H_2Cl_4"

"n(Cl_2) = 2n(C_2H_2Cl_4) = 2*7.50 = 15.00mol"

"V(Cl_2) = nV_m = 15.00*22.4 = 336L"